# Observing hidden neuronal states in experiments

### v2

@@ Abstract
Figure 3: VC and CC in silico. Protocol as described for Fig. 1 applied to (a) a type-I Morris-Lecar neuron model [30] as example of excitatory cell, and (b) a type-I Wang-Buzsaki neuron model [31] as example of inhibitory interneuron; see SI for parameter values. The two-dimensional fast subsystem has a S-shaped steady-state curve satisfying (4). The steady-state curve (4) and the (multi-color) S-shaped curve from the VC protocol indistinguishable throughout the range of input currents $I_{\mathrm{vc}}$. The orange curve resulting from the CC protocol is very close to $S^{0}$ and the VC protocol near its dynamically stable parts.
@@
Figure 4: VC and CC protocols applied to the Morris-Lecar model in a parameter regime where it behaves as a type-II neuron model.
@@
Geometrical singular perturbation theory (GSPT) by Fenichel [32] implies that after decay of initial transients every trajectory of the VC protocol modelled by system (1), (2) satisfies eq. (4) for the steady-state curve up to order $\varepsilon$. The first-order terms in $\varepsilon$ are
@@
$\begin{split}I_{\mathrm{eq}}(V)-I_{\mathrm{vc}}(V)&\approx I_{\mathrm{eq}}^{\prime}(V)\dot{V}\,\frac{\tau_{1/2}}{\ln 2}\\ &\lesssim I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})0.2\mbox{\,mV},\end{split}$ (5)
@@
where $\tau_{1/2}$ is the time for deviations from the transversally stable steady-state curve (4) to decay to half of their initial value; see SI for details (eq. (S7) and Fig. S3). We estimate $\tau_{1/2}$ from recovery transients after disturbances naturally occuring from imperfections in the voltage clamp during VC runs as $\tau_{1/2}\lesssim 0.075$ s (see SI, Fig. S3(b)), such that $\dot{V}\tau_{1/2}/\ln 2\approx\dot{V}_{\mathrm{h}}\tau_{1/2}/\ln 2\lesssim 0.2$ mV. Thus, the systematic bias between $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})$ in Fig. 1(b) and the true steady state curve $I_{\mathrm{eq}}(V)$ caused by dynamically changing $V_{\mathrm{h}}$ is below measurement disturbances.
@@
Estimate (5) can also be tested in silico. Figures 3 and 4 emulate both VC and CC protocols with the Morris-Lecar model [30], a biophysical model typically used for excitatory neurons of either type-I (Fig. 3(a)) or type-II (Fig. 4) excitability, and with the Wang-Buzsaki model (Fig. 3(b)), which is a classical model of inhibitory interneuron. Both models are of form (1) ($j=n_{j}=1$). For the chosen parameter set (see SI), the curves $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ and $(I_{\mathrm{eq}}(V_{\mathrm{h}}(t)),V_{\mathrm{h}}(t))$ are order $\varepsilon$ ($\approx 1$%) apart.
@@
Consequently, time profile $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ follows closely the curve (4) of stable steady states of the fast subsystem (1) with VC protocol $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, treating $V_{\mathrm{h}}$ as a parameter. We now connect the curve (4) to a curve of fast-subsystem equilibria of the CC protocol, which is in part unstable. To this end we recast the VC protocol in the form of a CC protocol with non-constant current ramp speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$ and disturbances: the smoothed time profile $I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))$ in Fig. 1(b) of the VC run (thick, in blue/brown/red) equals the raw-data measured time profile $I_{\mathrm{vc}}(t)$ (thin blue curve with fluctuations) plus disturbances $\eta_{\mathrm{vc}}(t)$, defined by $\eta_{\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))-I_{\mathrm{vc}}(t)$. After smoothing, the derivative $I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}})$ w.r.t. $V_{\mathrm{h}}$ is moderate ($\lesssim 20$ pA///mV in modulus at its maximum near $I_{\mathrm{db}}$), such that the VC protocol implies
@@
where $\Delta_{I,\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$, hence $\Delta_{I,\mathrm{vc}}(t)\approx I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$, with upper bound $\max_{t}|\Delta_{I,\mathrm{vc}}(t)|\lesssim 3.7$ pA///ms in the range of Fig. 1(b). Thus, $I_{\mathrm{vc,sm}}$ is indeed still slow. Hence, except for disturbances $\eta_{\mathrm{vc}}(t)$, the external current $I_{\mathrm{ext}}$ is slowly varying according to a CC protocol with slowly time-varying speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$, such that the VC protocol (2) is equivalent to the CC protocol (6) with disturbances $\eta_{\mathrm{vc}}$.
Systems (1), (3) and (1), (6) without disturbances ($\eta_{\mathrm{vc}}=0$) are both models of CC protocols. They have the same fast subsystem (1) when setting $\varepsilon=0$ (i.e., for constant input current) and identifying $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$. The respective fast-subsystem steady states $(V,x_{j})_{j=1,\ldots,N}$ satisfy
@@
However, systems (1), (3) and (1), (6) differ by the nature of their respective slow variables $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$: $I_{\mathrm{h}}$ is an externally applied hold current for (3), while $I_{\mathrm{vc,sm}}$ is a measured (and smoothed) current from the feedback control $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ of the voltage source for (6). Thus, while the S-shaped steady-state curve $(I_{\mathrm{eq}}(V),V)$ is identical for both systems, it contains large unstable segments as a steady-state curve of (1), (3), while it is always stable as a steady-state curve of (1), (6). The change in stability is caused by the disturbances $\eta_{\mathrm{vc}}$, which are current adjustments generated by the feedback term in (2), $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. Along most of the curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ the $\eta_{\mathrm{vc}}$ are small fluctuations such that $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})\approx I_{\mathrm{vc}}(V_{\mathrm{h}})$ and the feedback is approximately non-invasive [7]. Estimate (5) ensures that the measurements $I_{\mathrm{vc}}(V)$ stay close to $I_{\mathrm{eq}}(V)$. Therefore, we can conclude that the VC protocol (2) with slowly varying feedback reference signal $V_{\mathrm{h}}$ reveals the entire family of steady states of a neuron (type 1 or 2) with constant external current $I_{\mathrm{ext}}$, both stable (observable) and unstable (non-observable, hidden). Consequently, the N-shaped $I$-$V$ relations for type-1 neurons reported in [19, 20, 23] equal S-shaped steady-state bifurcation diagrams for these neurons with respect to $I_{\mathrm{ext}}$. They are tractable with a VC protocol where the current $I_{\mathrm{ext}}$ is a sufficiently slowly varying feedback current with sufficiently small fluctuations $\eta_{\mathrm{vc}}$. In particular, this allows us to detect and pass through fold bifurcations directly in the experiment.
Figure 5: Periodic spiking responses to step-current protocols for PY cells 5 and 1. Panels (a1-b1) are reproductions of Fig. 1(a) of the main manuscript and of Fig. S1(a), respectively, with a dashed rectangle chosen around $I_{\mathrm{h}}=200$ pA. The corresponding zoomed view is provided in panels (a2-b2). Panels (a3-b3) show the time profiles of voltage responses $V_{\mathrm{cc}}$ from a current-clamp stimulation with constant hold current $I_{\mathrm{h}}=200$ pA; the red dashed lines highlight the transient convergence towards a periodic response, which we use to compute $\tau_{1/2}$. Finally, panels (a4-b4) show a phase-plane projection of the time series shown in panels (a3-b3), namely plotting $V_{\mathrm{cc}}^{\prime}$ against $V_{\mathrm{cc}}$. This shows that the responses stabilise onto a periodic solution.
@@
In contrast, for the CC open-loop protocol (3) applying a slowly varying electrical current the neuron dynamically transitions between its observable rest states and its observable (dynamically stable) spiking states (see Fig. 1(b) orange time profile). The speed of variation $\varepsilon\Delta_{I}\lesssim 0.75\times 10^{-3}$ pA///ms is such that $I_{\mathrm{ext}}$ varies by only about $1$ pA per spiking period. Transients to step current responses in the stable spiking region have a half-time for decay $\tau_{1/2}\lesssim 0.2$ s (see SI, Fig. S5 for the response to a step to $I_{\mathrm{ext}}=200$ pA), permitting us to estimate the systematic bias between true steady-state spiking and the response to dynamically changing $I_{\mathrm{ext}}$, similar to (5) for certain features of the periodic spike train, as described below.
As the description of the protocol in the main manuscript states we stimulate the neuron under current clamp with a current step during a calibration phase. This is done for a number of values of the step current. Panels (a2) and (b2) in Fig. 5 show the time profiles of the recorded voltage $V_{\mathrm{cc}}$ for a current step to $I_{\mathrm{h}}=200$ pA for PY cells 5 and 1. Panels (a1) and (a2) show where the end point of the step ($I_{\mathrm{h}}=200$ pA) is located in the respective bifurcation diagrams (vertical dashed line) for PY cells 5 and 1, respectively. The two dots along the vertical dashed line at $I_{\mathrm{h}}=200$ pA in panels (a1) and (a2) mark the maximum and minimum values of the voltage response (after transient) to the step-current stimulation, as shown in panels (a3) and (b3); panels (a2) and (b2) are zoomed views of panels (a1) and (b1), respectively, near $I_{\mathrm{h}}=200$, where the current-clamp data with constant hold current are considered. We observe that both responses to this step-current stimulation lead to a periodic spiking behaviour, after initial transients have decayed. This is especially visible on the bottom plots, obtained by approximating the time derivative of the voltage time series, which is plotted against the voltage, hence giving a phase diagram. In both case, the phase-plane trajectory reveals a convergence towards a periodic (spiking) state (inner oscillations). The transient (outer oscillations) seems to indicate a weak stability of these spiking states; a blue line indicates the direction of time, with two dots corresponding to points at times $t_{n}$ and $t_{n+m}$.
The black dots in panel (a1) indicate that for PY cell 5 there is quantitative agreement between the response to the CC protocol with slowly-varying applied current $I_{\mathrm{h}}(t)$ and the response to the current step. There is less agreement for PY cell 1, between step response and the response to the CC protocol with slow variation of $I_{\mathrm{h}}$. The time profile suggests that there are either different possible spiking responses and, hence, more bifurcations in the dynamic protocol, such as period-doubling bifurcations, or the condition of the PY cell has drifted between step response calibration and CC ramp.
We use the decay rate of transients toward the stable spiking response in panels (a3) and (b3) to estimate time to half-decay of transients $\tau_{1/2}$ for spiking responses in the stable part of the spiking region (similar to the estimate for $\tau_{1/2}$ from the large-fluctuation decay in Fig. S3 in Section S3 for eq. (5)).
@@
Observing how the spike maxima settle to their steady-state value in Fig. 5(a3), we conservatively estimate the half-time for the decay to half the amplitude deviation from steady-state maximum as $\tau_{1/2}\leq 0.2$ s. This half-time enters estimates for the bias caused by varying the hold current $I_{\mathrm{h}}$ dynamically with speed $\varepsilon\Delta_{I}$. For example, the estimate for the bias in the minimum voltage $V_{\min}$ of the spike equals to first order in $\varepsilon$ is
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}.$
@@
We can see in Fig. 5(a) that $V_{\min}(I_{\mathrm{h}})$ changes in the spiking region with about $8$ mV per $120$ pA, so $V_{\mathrm{min}}^{\prime}(I)\approx 0.07$ mV/\!//pA. The current changes with $\mathrm{d}I/\mathrm{d}t\approx 7$ pA///s. Thus, to first order in $\varepsilon$, the error then amounts to:
@@
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}\approx 0.14\mbox{\,mV.}$.
@@
Hence, the effect from changing the current $I_{\mathrm{h}}$ dynamically is small. The bias estimate grows to infinity when the amplitude $V_{\max}(I_{\mathrm{h}})$ envelope assumes a square-root like shape and $\tau_{1/2}$ goes to infinity, as is the case at the Hopf bifurcation.
@@
Thus, combining the VC protocol (6) for varying $I_{\mathrm{vc,sm}}$, and the CC protocol (3) for varying $I_{\mathrm{h}}$, enables us to interpret the data sets from both protocols in Fig. 1(b) as a bifurcation diagram including unstable states.
@@
For the experimental curves presented in Fig. 1(b,c) (and also in Fig. S1), the disturbances $\eta_{\mathrm{vc}}$ are not small in some unstable parts of the reported steady-state curve (e.g., near $V_{\mathrm{h}}=-30$ mV in Fig. 1(b)), caused by imperfect voltage clamping across the membrane (see Fig. SI3(b) in SI for a zoom into these intermittent dropout events). Furthermore, the distances between $(I_{\mathrm{vc,sm}}(t),V_{\mathrm{h}}(t))$ in the VC run and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ in the CC run for the same cell are visibly larger along parts of the curve corresponding to dynamically stable stationary states. This is due to the natural drift of the neuron’s physiological properties as it changes conductance properties (e.g., temperature, degradation,..) over time.
Conclusion.–

### v3

@@ 2.1 Analysis
Figure 3: VC and CC in silico. Protocol as described for Fig 1B applied to A: a class-I Morris-Lecar neuron model (10) [morris1981] as example of excitatory cell, and B: a class-I Wang-Buzsáki neuron model (12) [wang1996] as example of inhibitory interneuron; see (10) and (12) for differential equations and Tables 1 and 2 for parameter values. The two-dimensional fast subsystem has a S-shaped steady-state curve satisfying steady-state conditions (4). The steady-state IIVV curve (4) and the (multi-color) S-shaped curve from the VC protocol (2) are indistinguishable throughout the range of input currents $I_{\mathrm{vc}}$. The orange curve resulting from the CC protocol is very close to $S^{0}$ and the VC protocol near its dynamically stable parts. See Figs F and G in S1 Text for numerical bifurcation diagrams.
@@
Figure 4: VC and CC protocols as described for Fig 1B applied to the Morris-Lecar model in a parameter regime where it behaves as a class-II neuron model. See (10) for differential equations and Table 1 for parameter values and Fig H in S1 Text for numerical bifurcation diagram.
@@
Geometrical singular perturbation theory (GSPT) by Fenichel [fenichel1979] implies that after decay of initial transients every trajectory of the general model (1) with VC protocol (2) satisfies the algebraic conditions (4) for the steady-state curve up to order $\varepsilon$. The first-order terms in $\varepsilon$ are
@@
$\begin{split}I_{\mathrm{eq}}(V)-I_{\mathrm{vc}}(V)&\approx I_{\mathrm{eq}}^{\prime}(V)\dot{V}\,\frac{\tau_{1/2}}{\ln 2}\lesssim I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}})0.2\mbox{\,mV},\end{split}$ (5)
@@
where $\tau_{1/2}$ is the time for deviations from the transversally stable steady-state IVI-V curve (4) to decay to half of their initial value; see eq (7) and Fig C in S1 Text for details. We estimate $\tau_{1/2}$ from recovery transients after disturbances naturally occuring from imperfections in the voltage clamp during VC runs as $\tau_{1/2}\lesssim 0.075$ s (see Fig CB in S1 Text), such that $\dot{V}\tau_{1/2}/\ln 2\approx\dot{V}_{\mathrm{h}}\tau_{1/2}/\ln 2\lesssim 0.2$ mV. Thus, the systematic bias between $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})$ in Fig 1B and the true steady state curve $I_{\mathrm{eq}}(V)$ caused by dynamically changing $V_{\mathrm{h}}$ is below measurement disturbances.
@@
Deviation estimate (5) can also be tested in silico. Figs 3 and 4 emulate both VC and CC protocols with the Morris-Lecar model [morris1981] (see Materials and Methods, equation (10)), a biophysical model typically used for excitatory neurons of either class-I (Fig 3A) or class-II (Fig 4) excitability, and with the Wang-Buzsáki model (Fig 3B), which is a classical model of inhibitory interneuron [wang1996] (see Materials and Methods, equation (12)). Both models are of general form (1) with $j=n_{j}=1$. For the chosen parameter set (see S1 Text), the curves $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ and $(I_{\mathrm{eq}}(V_{\mathrm{h}}(t)),V_{\mathrm{h}}(t))$ are order $\varepsilon$ ($\approx 1$%) apart.
Consequently, time profile $(I_{\mathrm{vc}}(t),V_{\mathrm{h}}(t))$ follows closely the steady-state IVI-V curve (4) of stable steady states of the fast subsystem (1) with VC protocol $I_{\mathrm{ext}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$, treating $V_{\mathrm{h}}$ as a parameter. This implies that faster ramp speeds are permissible when optimising trade-off between drift of cell properties and bias due to nonzero ramp speed. At high voltages the factor $I_{\mathrm{eq}}^{\prime}(V)$ becomes large, such that estimate (5) predicts larger deviations for large $V_{\mathrm{h}}$, as confirmed in Figs 1B, 2, AB, AC and AD in S1 Text.
@@
We now connect the steady-state IVI-V curve (4) of the VC protocol to a curve of fast-subsystem equilibria of the CC protocol, which is in part unstable. To this end we recast the VC protocol in the form of a CC protocol with non-constant current ramp speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$ and disturbances: the smoothed time profile $I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))$ in Fig 1B of the VC run (thick, in blue/brown/red) equals the raw-data measured time profile $I_{\mathrm{vc}}(t)$ (thin blue curve with fluctuations) plus disturbances $\eta_{\mathrm{vc}}(t)$, defined by $\eta_{\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}(V_{\mathrm{h}}(t))-I_{\mathrm{vc}}(t)$. After smoothing, the derivative $I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}})$ w.r.t. $V_{\mathrm{h}}$ is moderate ($\lesssim 20$ pA//mV in modulus at its maximum near $I_{\mathrm{db}}$), such that during the VC protocol the external current $I_{\mathrm{ext}}$ satisfies
@@
where $\Delta_{I,\mathrm{vc}}(t)=I_{\mathrm{vc,sm}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$. Hence, $\Delta_{I,\mathrm{vc}}(t)\approx I_{\mathrm{eq}}^{\prime}(V_{\mathrm{h}}(t))\Delta_{V}$, with upper bound $\max_{t}|\Delta_{I,\mathrm{vc}}(t)|\lesssim 3.7$ pA//ms in the range of Fig 1B. Thus, $I_{\mathrm{vc,sm}}$ is indeed still slow. So, except for disturbances $\eta_{\mathrm{vc}}(t)$, the external current $I_{\mathrm{ext}}$ is slowly varying according to a CC protocol with slowly time-varying speed $\varepsilon\Delta_{I,\mathrm{vc}}(t)$, such that the VC protocol (2) is equivalent to the CC protocol (6) with disturbances $\eta_{\mathrm{vc}}$.
This means that general model (1) with current $I_{\mathrm{ext}}$ given in (6) with zero disturbances ($\eta_{\mathrm{vc}}=0$) is a model for a CC protocol with driving current $I_{\mathrm{vc,sm}}$, in contrast to the model with open-loop CC protocol (1), (3). Both models have the same fast subsystem (1) when setting $\varepsilon=0$ (i.e., for constant input current) and identifying $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$. The respective fast-subsystem steady states $(V,x_{j})_{j=1,\ldots,N}$ satisfy
@@
However, the two models differ by the nature of their respective slow variables $I_{\mathrm{h}}$ and $I_{\mathrm{vc,sm}}$: $I_{\mathrm{h}}$ is an externally applied hold current for the open-loop CC protocol (3), while $I_{\mathrm{vc,sm}}$ is a measured (and smoothed) current from the closed-loop feedback control $g_{\mathrm{c}}(V_{\mathrm{h}}-V)$ of the voltage source for (6). Thus, while the S-shaped steady-state curve $(I_{\mathrm{eq}}(V),V)$ is identical for both models, it contains large unstable segments as a steady-state curve of open-loop CC protocol (1), (3), while it is always stable as a steady-state curve of closed-loop VC protocol (1), (6). The change in stability is caused by the disturbances $\eta_{\mathrm{vc}}$, which are current adjustments generated by the feedback term in VC protocol (2), $I_{\mathrm{vc}}=g_{\mathrm{c}}(V_{\mathrm{h}}-V)$. Along most of the curve $(I_{\mathrm{vc,sm}}(V_{\mathrm{h}}),V_{\mathrm{h}})$ the $\eta_{\mathrm{vc}}$ are small fluctuations such that $I_{\mathrm{vc,sm}}(V_{\mathrm{h}})\approx I_{\mathrm{vc}}(V_{\mathrm{h}})$ and the feedback is approximately non-invasive [sieber2008a]. Estimate (5) ensures that the measurements $I_{\mathrm{vc}}(V)$ stay close to $I_{\mathrm{eq}}(V)$. Therefore, we can conclude that the VC protocol (2) with slowly varying feedback reference signal $V_{\mathrm{h}}$ reveals the entire family of steady states of a neuron (class I or II) with constant external current $I_{\mathrm{ext}}$, both stable (observable) and unstable (non-observable, hidden). Consequently, the N-shaped II-$V$ relations for class-I neurons reported in [schwindt1981, fishman1969, vervaeke2006] equal S-shaped steady-state bifurcation diagrams for these neurons with respect to $I_{\mathrm{ext}}$. They are tractable with a VC protocol where the current $I_{\mathrm{ext}}$ is a sufficiently slowly varying feedback current with sufficiently small fluctuations $\eta_{\mathrm{vc}}$. In particular, this allows us to detect and pass through fold bifurcations directly in the experiment.
Figure 5: Periodic spiking responses to step-current protocols for PY cells 5 (left column) and 1 (right column). Panels A, B: time profiles of voltage responses $V_{\mathrm{cc}}(t)$ from a current-clamp stimulation with a step to constant hold current $I_{\mathrm{h}}=200$ pA from $0.1$ to $1.6$ s (0.10.10.60.6 s in blue, 0.60.61.61.6 s in black). Red markings show how half-decay time $\tau_{1/2}$ is extracted from voltage maxima during transients. Panels C, D: $(V_{\mathrm{cc}},V_{\mathrm{cc}}^{\prime})$ phase-plane projection the time series from panels A, B, using a one-step finite-difference approximation of $V_{\mathrm{cc}}^{\prime}(t)$ with color code matching panels A, B to distinguish transients and steady-state spiking. Panels E, F: reproduction of Figs 1A and $A$ in S1 Text without bifurcation or stability markings, respectively. The value $I_{\mathrm{h}}=200$ pA (vertical dotted black line) and a $10$ ms window around it (vertical solid black lines) are highlighted. The insets show a zoom into this $10$ ms window around $I_{\mathrm{h}}=200$ pA of panels E, F. Black crosses in panels E, F are minimum and maximum of steady-state spiking from panels A, B for comparison.
@@
In contrast, during the CC open-loop protocol (3), when applying a slowly varying electrical current, the neuron dynamically transitions between its observable rest states and its observable (dynamically stable) spiking states (see Fig 1B orange time profile). The speed of variation $\varepsilon\Delta_{I}\lesssim 0.75\times 10^{-3}$ pA//ms is such that $I_{\mathrm{ext}}$ varies by $1$ pA or less per spiking period. For a transient decay analysis we stimulate the neuron with larger current steps during a calibration phase before executing CC protocol runs. Fig 5A and 5B show the response to such a current step to $I_{\mathrm{ext}}=200$ pA for cells $5$ and $1$. We observe that the transients in the step current response such as in Fig 5A and 5B in the stable spiking region have a half-time for decay toward the stable spikes
$\tau_{1/2}\lesssim 0.2\mbox{\,s.}$ (8)
@@
This half-time $\tau_{1/2}$ enters estimates for the bias caused by varying the hold current $I_{\mathrm{h}}$ dynamically with speed $\varepsilon\Delta_{I}$. For example, the estimate for the bias in the minimum voltage $V_{\min}$ of the spike equals to first order in $\varepsilon$
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}.$ (9)
@@
We can see in Fig 5E and 5F that $V_{\min}(I_{\mathrm{h}})$ changes in the spiking region with about $8$ mV per $120$ pA, so $V_{\mathrm{min}}^{\prime}(I)\approx 0.07$ mV/\!/pA. The current changes with $\mathrm{d}I/\mathrm{d}t\approx 7.5$ pA//s. Thus, to first order in $\varepsilon$, the error then amounts to:
@@
$\operatorname{bias}(V_{\min})\approx V_{\min}^{\prime}(I_{\mathrm{h}})\times\frac{\mathrm{d}}{\mathrm{d}t}I_{\mathrm{h}}(t)\times\frac{\tau_{1/2}}{\ln 2}\approx 0.14\mbox{\,mV.}$
@@
Hence, the effect from changing the current $I_{\mathrm{h}}$ dynamically is small (below visibility in bifurcation diagrams such as Fig 1B or Fig 5E and 5F. The linear bias estimate grows to infinity when the amplitude $V_{\max}(I_{\mathrm{h}})$ envelope assumes a square-root like shape and $\tau_{1/2}$ goes to infinity, as is the case at Hopf bifurcations.
@@
Thus, combining the VC protocol (6) for varying $I_{\mathrm{vc,sm}}$, and the CC protocol (3) for varying $I_{\mathrm{h}}$, enables us to interpret the data sets from both protocols in Fig 1B as a bifurcation diagram including unstable states.
Drift and intermittent dropouts
In Fig 5E and 5F the black markers (++) at $I_{\mathrm{h}}=200$ pA mark the maximum and minimum values of the voltage response (after transient) to the step-current stimulation, as shown in panels A and B. They indicate that for PY cell 5 there is quantitative agreement between the response to the CC protocol with slowly-varying applied current $I_{\mathrm{h}}(t)$ and the response to the current step. There is less agreement for PY cell 1, between step response and the response to the CC protocol with slow variation of $I_{\mathrm{h}}$. The time profile suggests that there are either different possible spiking responses and, hence, more bifurcations in the dynamic protocol, such as period-doubling bifurcations, or the condition of the PY cell has drifted between step response calibration and CC ramp.
@@
Furthermore, the distances between $(I_{\mathrm{vc,sm}}(t),V_{\mathrm{h}}(t))$ in the VC run and $(I_{\mathrm{h}}(t),V_{\mathrm{cc}}(t))$ in the CC run for the same cell are visibly larger than predicted by eq (5) along parts of the curve corresponding to dynamically stable stationary states. This is due to the natural drift of the neuron’s physiological properties as it changes conductance properties (e.g., temperature, degradation,…). In S1 Text we show that over time.
Finally, for the experimental curves presented in Figs 1B, 1C, 2 and A in S1 Text, the disturbances $\eta_{\mathrm{vc}}$ are not small in some unstable parts of the reported steady-state curve (e.g., near $V_{\mathrm{h}}=-30$ mV in Fig 1B), caused by imperfect voltage clamping across the membrane. Fig CB in S1 Text shows a zoom demonstrating that the observed current spikes are indeed intermittent dropout events. Detailed simulations using Morris-Lecar model (10), shown in S1 Text Section 3, reproduce these dropouts closely, if adds short current spikes and small-amplitude white noise to (10).
3 Discussion